At a sharp horizontal circular turn a bus slows down going from 90.0 km/h to 50.0 km/h in the 15.0 s it takes to round the bend. The radius of the curve is 150 m. Assume the bus continues to slow down at the same rate. At the moment the bus speed reaches 50.0 km/h calculate: (i)the magnitude and direction of the tangential acceleration. (ii)the radial acceleration (magnitude and direction). (iii)the magnitude and direction of the total acceleration


1
Expert's answer
2021-12-10T11:31:34-0500
"a_t=\frac{13.9-25}{15}=-0.741\frac{m}{s^2}\\
a_n=\frac{13.9^2}{150}=1.29\frac{m}{s^2}\\\theta=29.9\degree\\a=1.48\frac{m}{s^2}"

inward and 29.90 backward.